# Identification and estimation of mediational effects of longitudinal modified treatment policies

### v3

@@ Abstract
We demonstrate a comprehensive semiparametric approach to causal mediation analysis, addressing the complexities inherent in settings with longitudinal and continuous treatments, confounders, and mediators. Our methodology utilizes a nonparametric structural equation model and a cross-fitted sequential regression technique based on doubly robust pseudo-outcomes, yielding an efficient, asymptotically normal estimator without relying on restrictive parametric modeling assumptions. We are motivated by a recent scientific controversy regarding the effects of invasive mechanical ventilation (IMV) on the survival of COVID-19 patients, considering acute kidney injury (AKI) as a mediating factor. We highlight the possibility of “inconsistent mediation,” in which the direct and indirect effects of the exposure operate in opposite directions. We discuss the significance of mediation analysis for scientific understanding and its potential utility in treatment decisions.mediation, machine learning, longitudinal, treatment policies
@@ 1 Introduction
Causal mediation analysis, or the identification of the different pathways through which an exposure or treatment (we will use the terms “exposure” and “treatment” interchangeably) can operate, has been increasing in interest across the scientific literature (Nguyen et al., 2021). Although the concept of mediation has been active for at least a century (Wright, 1921), mediation analysis has only relatively recently been incorporated in formal causal models such as the framework of potential outcomes or structural causal models, which allow for nonparametric perspectives, or the study of mediation effects without reference to a particular (e.g., linear) statistical model (Imai et al., 2010).
@@
Mediation analysis in longitudinal settings is particularly challenging; some existing approaches have relied on parametric model specification (VanderWeele and Tchetgen, 2017; Tai et al., 2022), which leads to a lack of robustness. Zheng and van der Laan (2017) provided a robust machine-learning approach for the estimation of mediation effects with time-varying treatments, confounders, and mediators. This line of research was advanced, for example, by Díaz et al. (2023b) which allowed the (interventional) direct effect to include pathways through intermediate confounders, which allows the indirect effect to isolate precisely those operating through the mediator and not other variables.
@@
However, even these strategies that make use of flexible models for estimation still rely on the notion of a “static” treatment regimen, or one in which each value of treatment for all timepoints is explicitly specified in the estimand of interest, which is a significant limitation insofar as real-world decision-making rarely follows this form in the longitudinal setting. That is, even an externally-designed intervention is likely to be responsive, intentionally or otherwise, to variables that are encountered over the course of treatment. A few examples include the administration of antiretroviral drugs in HIV patients, which can be adapted to CD4 cell count measurements (Hernán et al., 2006), and the administration of antihypertensive drugs, which can be adapted to blood pressure measurements (Johnson et al., 2018).
@@
Aside from their lack of applicability, studies of static intervention face various statistical challenges. First, identification of the causal estimand relies on the positivity assumption (i.e., that there is a positive probability of any treatment regime given any level of covariates), which is often implausible, especially for continuous treatments given at multiple timepoints. Likewise, for continuous treatments, the estimation of a dose-response curve (the function of exposure values that yields the population-average outcome conditional on receiving each value) involves complications in terms of convergence rates (standard $\sqrt{n}$ rates are not possible, see Kennedy, 2018) and of the summarization and interpretation of such a high-dimensional target parameter.
@@
By contrast, so-called “dynamic” treatment regimes allow exposure values to change based on observed covariates (Robins et al., 2004). These can be generalized to “stochastic” interventions which allow for randomness in the exposure values, even conditional on all other variables. Furthermore, it may often be of interest to consider a treatment regime that assigns treatment values depending on what treatment value would have occurred in the absence of intervention, or the “natural” value of treatment (Young et al., 2014). Such a regime, which we call a “modified treatment policy,” was considered in Robins et al. (2004), with formalization and estimation strategies for single-timepoint studies to follow in Díaz and van der Laan (2012) and Haneuse and Rotnitzky (2013). Díaz et al. (2021) provide a framework for robust estimation of the effects of modified treatment policies for time-varying treatments.
@@
We synthesize these various strands of the statistical literature into a new strategy for mediation analysis in longitudinal settings, following Díaz and Hejazi (2020) and Hejazi et al. (2023), who studied mediation analysis for single-timepoint stochastic interventions. Specifically, we consider the direct and indirect effects of longitudinal modified treatment policies under interventional mediator distributions, which have not been considered in previous literature. The methods we propose are motivated by an issue that was relevant in the early landscape of COVID-19 management: the use of invasive mechanical ventilation (IMV) in cases of acute respiratory distress syndrome (ARDS). ARDS is a form of acute hypoxic respiratory failure, which is one of the key time-varying features of a patient’s course with severe COVID (see, e.g., Wang et al. (2021)). In the beginning stages of the pandemic, IMV was employed early, or at a lower supplemental oxygen delivery threshold, to manage ARDS, but its usage has been associated with several iatrogenic risks such as ventilator-associated pneumonia (Wicky et al., 2021) and barotrauma (Shrestha et al., 2022). Of interest to the present study is acute kidney injury (AKI), a critical condition that complicates ICU stays and is associated with mortality (Vemuri et al., 2022). As more data emerged and practitioners became more familiar with the pace of disease progression, guidelines shifted towards delaying intubation due to high mortality rates among mechanically ventilated patients and potential secondary complications. To navigate these treatment dynamics, we aim to elucidate the direct and indirect effects of IMV on patient mortality with respect to the AKI pathway. This application offers a compelling case for the use of our methods in understanding the effects of time-varying treatments and the causal pathways involved. We emphasize the phenomenon of “inconsistent mediation” (a situation in which the direct and indirect effects are qualitatively in opposite directions). In the current setting, this would involve (on average) IMV causing death through AKI while (on average) preventing death through other mechanisms. This dichotomy may be relevant for research and clinical practice, if interventions could be in place to mitigate IMV-induced AKI. While this is an active area of research, potential interventions include the administration of diuretics to modulate intravascular volume status (Grams et al., 2011; Glassford and Bellomo, 2011) and pharmacological options as explored by Pickkers et al. (2022). Finally, we also investigate the relationship between baseline covariates and the magnitude of these direct and indirect effects; such relationships can suggest biological pathways as well as potentially inform future treatments.
@@ 2.1 Notation
Our notation largely mirrors that of Díaz et al. (2021) and Díaz et al. (2023b). We consider data $X_{1},\ldots,X_{n}$, an i.i.d. sample from a distribution $\mathsf{P}$. The data are longitudinal, with $X=(L_{1},A_{1},Z_{1},M_{1},L_{2},\ldots,A_{\tau},Z_{\tau},\allowbreak M_{\tau},L_{\tau+1})$, where the variables $A$ are the exposure of interest, $L$ are covariates, $M$ are mediators, and $Z$ are intermediate confounders. An intermediate confounder is a variable affected by exposure and which affects the mediator and outcome; see VanderWeele et al. (2014). We let $Y=L_{\tau+1}$ denote the outcome of interest. In our illustrative application, $A$ will denote intubation status, $M$ is the occurence of AKI, $Y$ is mortality/survival, and $L$ and $Z$ are covariates or lab results, either at baseline or during the hospital stay. As in our illustrative application, we will assume that the mediator takes values in a finite set; the supports of all other variables are unrestricted.
@@ 2.2 Structural equation model and definition of causal effects
We consider causal effects defined within a nonparametric structural equation model, which consists of exogenous and endogenous variables and unknown deterministic functions governing the causal relations between them, together with independence assumptions on the exogenous variables (Pearl, 2000). We allow for arbitrary causal relations among observed variables that respect their time-ordering; this can be visualized in the saturated directed acyclic graph (DAG) given in Figure 1, reproduced from Díaz et al. (2023b).
@@
Figure 1: Causal DAG for the model of time-varying mediation. For clarity, the symbol $\boldsymbol{\mapsto}$ is used to indicate arrows from all previous nodes (to the left) to all following nodes (to the right).
@@
This setup allows us to speak of causal effects as the interventions on some functions $f$ while keeping the remaining structure of the model intact. For example, the simple intervention to set $A_{1}$ to the value $a_{1}$ replaces $f_{A,1}$ with the constant function $a_{1}$. In this paper, we are interested in the effects of interventions on the treatment variables $\bar{A}$, and we will analyze its decomposition through pathways that include mediating variables $\bar{M}$ and through pathways that do not. For this purpose, we will consider interventions on the functions $f_{A,t}$ and $f_{M,t}$ for $t=1,\ldots,\tau$.
For interventions on the treatment variable, we will consider longitudinal modified treatment policies (LMTPs) (Díaz et al., 2021), which are hypothetical interventions defined in terms of a sequence of functions $\mathsf{d}=(\mathsf{d}_{1},\ldots,\mathsf{d}_{\tau})$, where $\mathsf{d}_{t}$ is a function $\mathsf{d}_{t}(a_{t},h_{A,t},\epsilon_{t})$ that depends on a treatment value $a_{t}$, a history value $h_{A,t}$, and possibly a randomizer $\epsilon_{t}$. (We will not consider such a randomizer in this paper.) The intervention is defined by removing the equation $A_{t}=f_{A_{t}}(H_{A,t},U_{A,t})$ sequentially (from $t=1$ to $t=\tau$) from the structural model, and replacing it with evaluations of the functions $\mathsf{d}_{t}$ as follows. At the first timepoint, the LMTP assigns the exposure as a new random variable $A_{1}^{\mathsf{d}}=\mathsf{d}_{t}(A_{1},H_{A,1},\epsilon_{1})$. This generates a counterfactual treatment $A_{2}(\bar{A}_{1}^{\mathsf{d}})=f_{A_{2}}(H_{A,2}(\bar{A}_{1}^{\mathsf{d}}),U_{A,1})$ that would have been observed if the intervention had been discontinued right before $t=2$. Then, the intervention at $t=2$ assigns treatment as $A_{2}^{\mathsf{d}}=\mathsf{d}_{2}(A_{2}(\bar{A}_{1}^{\mathsf{d}}),H_{A,2}(\bar{A}_{1}^{\mathsf{d}}),\epsilon_{2})$.
Across multiple timepoints, it is helpful to define the counterfactual history $H_{A,t}(\bar{A}_{t-1}^{\mathsf{d}})$ as the sequence of variables prior to $A_{t}$ that would have been observed under the LMTP given by $\bar{A}_{t-1}^{\mathsf{d}}$. In addition, $A_{t}(\bar{A}_{t-1}^{\mathsf{d}})=f_{A_{t}}(H_{A,t}(\bar{A}_{t-1}^{\mathsf{d}}),U_{A,t})$, termed the “natural value of treatment” (Richardson and Robins, 2013; Young et al., 2014), “represents the value of treatment that would have been observed at time $t$ under an intervention carried out up until time $t-1$ but discontinued thereafter” (Díaz et al., 2021). The intervention at time $t$ is given by $A_{t}^{\mathsf{d}}=\mathsf{d}_{t}(A_{t}(\bar{A}_{t-1}^{\mathsf{d}}),H_{A,t}(\bar{A}_{t-1}^{\mathsf{d}}),\epsilon_{t})$.
@@ 3.1 Assumptions and sequential regression
To identify the parameter, $\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))]$, the following assumptions, which combine elements of the identification assumptions from Díaz et al. (2021) and Díaz et al. (2023b), are sufficient. Note that 1 involves future histories rather than single timepoints only.
@@
3Piecewise-smooth invertibililty of treatment modification.
@@
Our identification strategy makes use of a representation of longitudinal causal estimands by means of “sequential regression” as proposed by Bang and Robins (2005) and further investigated by van der Laan and Gruber (2012); Luedtke et al. (2017), and Rotnitzky et al. (2017). This approach considers a recursive regression of outcomes (or pseudo-outcomes) onto previous variables, starting at $t=\tau$ and continuing backward to $t=0$. As in Díaz et al. (2023b), we split the recursion over the two non-interventional variables $L$ and $Z$ as follows. We consider fixed, pre-specified interventions $\mathsf{d}^{\prime}$, $\mathsf{d}^{\star}$, and $\bar{m}$. Starting with, $\mathsf{Q}_{Z,\tau+1}=Y$, define
@@
When the meaning is clear, noting that the only non-random component of the functions $\mathsf{Q}$ are the user-specified $\mathsf{d}^{\prime},\mathsf{d}^{\star}$, and $\bar{m}$, we may use the abbreviations $\mathsf{Q}_{L,t}^{\prime}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{L,t}(\bar{H}_{M,t},\underaccent{\bar}{m}_{t})$, $\mathsf{Q}_{Z,t}^{\prime}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{Z,t}(A^{{}^{\prime}}_{t},\bar{H}_{A,t},\underaccent{\bar}{m}_{t})$, and $\mathsf{Q}_{M,t}^{\star}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{M,t}(A_{t}^{\star},\bar{H}_{A,t},\underaccent{\bar}{m}_{t})$. We give the following identification result, which is a generalization of Theorem 1 of Díaz et al. (2023b).

### v4

@@ Abstract
We demonstrate a comprehensive semiparametric approach to causal mediation analysis, addressing the complexities inherent in settings with longitudinal and continuous treatments, confounders, and mediators. Our methodology utilizes a nonparametric structural equation model and a cross-fitted sequential regression technique based on doubly robust pseudo-outcomes, yielding an efficient, asymptotically normal estimator without relying on restrictive parametric modeling assumptions. We are motivated by a recent scientific controversy regarding the effects of invasive mechanical ventilation (IMV) on the survival of COVID-19 patients, considering acute kidney injury (AKI) as a mediating factor. We highlight the possibility of “inconsistent mediation,” in which the direct and indirect effects of the exposure operate in opposite directions. We discuss the significance of mediation analysis for scientific understanding and its potential utility in treatment decisions.
@@ 1 Introduction
Causal mediation analysis, or the identification of the different pathways through which an exposure or treatment (we will use the terms “exposure” and “treatment” interchangeably) can operate, has seen increasing interest across the scientific literature (Nguyen et al., 2021). Although the concept of mediation has been active for at least a century (Wright, 1921), mediation analysis has only relatively recently been incorporated in formal causal models such as the framework of potential outcomes or structural causal models, which allow for nonparametric perspectives, or the study of mediation effects without reference to a particular (e.g., linear) statistical model (Imai et al., 2010).
@@
Mediation analysis in longitudinal settings is particularly challenging; some existing approaches have relied on parametric model specification (VanderWeele and Tchetgen, 2017; Tai et al., 2022), which leads to a lack of robustness. Zheng and van der Laan (2017) provided a robust machine-learning approach for the estimation of mediation effects with time-varying treatments, confounders, and mediators. This line of research was advanced, for example, by D´ıaz et al. (2023b) which allowed the (interventional) direct effect to include pathways through intermediate confounders, thereby allowing the indirect effect to isolate precisely those operating through the mediator and not other variables.
@@
However, even these strategies that make use of flexible models for estimation still rely on the notion of a “static” treatment regime, or one in which each value of treatment for all timepoints is explicitly specified in the estimand of interest, which is a significant limitation, given that real-world decision-making rarely follows this form in the longitudinal setting. That is, even an externally designed intervention is likely to be responsive, intentionally or otherwise, to variables that are encountered over the course of treatment. A few examples are the administration of antiretroviral drugs in HIV patients, which can be adapted to CD4 cell count measurements (Hernán et al., 2006), and the administration of antihypertensive drugs, which can be adapted to blood pressure measurements (Johnson et al., 2018).
@@
Aside from their lack of applicability, studies of static interventions face various statistical challenges. First, identification of the causal estimand relies on the positivity assumption (i.e., that there is a positive probability of any treatment course given any level of covariates), which is often implausible, especially for continuous treatments given at multiple timepoints. Likewise, for continuous treatments, the estimation of a dose-response curve (the function of exposure values that yields the population-average outcome conditional on receiving each value) involves complications in terms of convergence rates (standard $\sqrt{n}$ rates are not possible, see Kennedy, 2018) and of the summarization and interpretation of such a high-dimensional target parameter.
@@
By contrast, so-called “dynamic” treatment regimes allow exposure values to change based on observed covariates (Robins et al., 2004). These can be generalized to “stochastic” interventions which allow for randomness in the exposure values, even conditional on all other variables. Furthermore, it may often be of interest to consider a treatment regime that assigns treatment values depending on what treatment value would have occurred in the absence of intervention, or the “natural” value of treatment (Young et al., 2014). Such a regime, which we call a “modified treatment policy,” was considered in Robins et al. (2004), with formalization and estimation strategies for single-timepoint studies to follow in Díaz and van der Laan (2012) and Haneuse and Rotnitzky (2013). D´ıaz et al. (2021) provide a framework for robust estimation of the effects of modified treatment policies for time-varying treatments.
@@
We synthesize these various strands of the statistical literature into a new strategy for mediation analysis in longitudinal settings, following D´ıaz and Hejazi (2020) and Hejazi et al. (2023), who studied mediation analysis for single-timepoint stochastic interventions. Specifically, we consider the direct and indirect effects of longitudinal modified treatment policies under interventional mediator distributions, which have not been considered in previous literature. The methods we propose are motivated by an issue that was relevant in the early landscape of COVID-19 management: the use of invasive mechanical ventilation (IMV) in cases of acute respiratory distress syndrome (ARDS). ARDS is a form of acute hypoxic respiratory failure, which is one of the key time-varying features of a patient’s course with severe COVID (see, e.g., Wang et al. (2021)). In the beginning stages of the pandemic, IMV was employed early, or at a lower supplemental oxygen delivery threshold, to manage ARDS, but its usage has been associated with several iatrogenic risks such as ventilator-associated pneumonia (Wicky et al., 2021) and barotrauma (Shrestha et al., 2022). Of interest to the present study is acute kidney injury (AKI), a critical condition that complicates ICU stays and is associated with mortality (Vemuri et al., 2022). As more data emerged and practitioners became more familiar with the pace of disease progression, guidelines shifted towards delaying intubation due to high mortality rates among mechanically ventilated patients and potential secondary complications. To navigate these treatment dynamics, we aim to elucidate the direct and indirect effects of IMV on patient mortality with respect to the AKI pathway. This application offers a compelling case for the use of our methods in understanding the effects of time-varying treatments and the causal pathways involved. We emphasize the phenomenon of “inconsistent mediation” (a situation in which the direct and indirect effects are in opposite directions). In the current setting, this would involve (on average) IMV causing death through AKI while (on average) preventing death through other mechanisms. This dichotomy may be relevant for research and clinical practice, if interventions could be in place to mitigate IMV-induced AKI. While this is an active area of research, potential interventions include the administration of diuretics to modulate intravascular volume status (Grams et al., 2011; Glassford and Bellomo, 2011) and pharmacological options as explored by Pickkers et al. (2022). Finally, we also investigate the relationship between baseline covariates and the magnitude of these direct and indirect effects; such relationships can suggest biological pathways as well as potentially inform future treatments.
@@ 2.1 Notation
Our notation largely mirrors that of D´ıaz et al. (2021) and D´ıaz et al. (2023b). We consider data $X_{1},\ldots,X_{n}$, an i.i.d. sample from a distribution $\mathsf{P}$. The data are longitudinal, with $X=(L_{1},A_{1},Z_{1},M_{1},L_{2},\ldots,A_{\tau},Z_{\tau},\allowbreak M_{\tau},L_{\tau+1})$, where the variables $A$ are the exposure of interest, $L$ are covariates, $M$ are mediators, and $Z$ are intermediate confounders. Under right censoring, we let $A_{t}$ denote a bivariate variable that includes an indicator of censoring as well as the exposure, and let the data become degenerate for any timepoint after the censoring time, adopting the methodology described in D´ıaz et al. (2023a). An intermediate confounder is a variable affected by exposure and which affects the mediator and outcome; see VanderWeele et al. (2014). We let $Y=L_{\tau+1}$ denote the outcome of interest. In our illustrative application, $A$ will denote intubation status, $M$ is the occurrence of AKI, $Y$ is mortality/survival, and $L$ and $Z$ are covariates or lab results, either at baseline or during the hospital stay. As in our illustrative application, we will assume that the mediator takes values in a finite set; the supports of all other variables are unrestricted.
@@ 2.2 Structural equation model and definition of causal effects
We consider causal effects defined within a nonparametric structural equation model, which consists of exogenous and endogenous variables and unknown deterministic functions governing the causal relations between them, together with independence assumptions on the exogenous variables (Pearl, 2000). We allow for arbitrary causal relations among observed variables that respect their time-ordering; this can be visualized in the saturated directed acyclic graph (DAG) given in Figure 1, reproduced from D´ıaz et al. (2023b).
@@
[…]
@@
This setup allows us to speak of causal effects as the interventions on some functions $f$ while keeping the remaining structure of the model intact. For example, the simple intervention to set $A_{1}$ to the value $a_{1}$ replaces $f_{A,1}$ with the constant function $a_{1}$. In this paper, we are interested in the effects of interventions on the treatment variables $\bar{A}$, and we will analyze their decomposition through pathways that include mediating variables $\bar{M}$ and through pathways that do not. For this purpose, we will consider interventions on the functions $f_{A,t}$ and $f_{M,t}$ for $t=1,\ldots,\tau$.
For interventions on the treatment variable, we will consider longitudinal modified treatment policies (LMTPs) (D´ıaz et al., 2021), which are hypothetical interventions defined in terms of a sequence of functions $\mathsf{d}=(\mathsf{d}_{1},\ldots,\mathsf{d}_{\tau})$, where $\mathsf{d}_{t}$ is a function $\mathsf{d}_{t}(a_{t},h_{A,t},\epsilon_{t})$ that depends on a treatment value $a_{t}$, a history value $h_{A,t}$, and possibly a randomizer $\epsilon_{t}$. (We will not consider such a randomizer in this paper.) The intervention is defined by removing the equation $A_{t}=f_{A_{t}}(H_{A,t},U_{A,t})$ sequentially (from $t=1$ to $t=\tau$) from the structural model, and replacing it with evaluations of the functions $\mathsf{d}_{t}$ as follows. At the first timepoint, the LMTP assigns the exposure as a new random variable $A_{1}^{\mathsf{d}}=\mathsf{d}_{1}(A_{1},H_{A,1},\epsilon_{1})$. This generates a counterfactual treatment $A_{2}(\bar{A}_{1}^{\mathsf{d}})=f_{A_{2}}(H_{A,2}(\bar{A}_{1}^{\mathsf{d}}),U_{A,1})$ that would have been observed if the intervention had been discontinued right before $t=2$. Then, the intervention at $t=2$ assigns treatment as $A_{2}^{\mathsf{d}}=\mathsf{d}_{2}(A_{2}(\bar{A}_{1}^{\mathsf{d}}),H_{A,2}(\bar{A}_{1}^{\mathsf{d}}),\epsilon_{2})$.
Across multiple timepoints, it is helpful to define the counterfactual history $H_{A,t}(\bar{A}_{t-1}^{\mathsf{d}})$ as the sequence of variables prior to $A_{t}$ that would have been observed under the LMTP given by $\bar{A}_{t-1}^{\mathsf{d}}$. In addition, $A_{t}(\bar{A}_{t-1}^{\mathsf{d}})=f_{A_{t}}(H_{A,t}(\bar{A}_{t-1}^{\mathsf{d}}),U_{A,t})$, termed the “natural value of treatment” (Richardson and Robins, 2013; Young et al., 2014), “represents the value of treatment that would have been observed at time $t$ under an intervention carried out up until time t1t-1 but discontinued thereafter” (D´ıaz et al., 2021). The intervention at time $t$ is given by $A_{t}^{\mathsf{d}}=\mathsf{d}_{t}(A_{t}(\bar{A}_{t-1}^{\mathsf{d}}),H_{A,t}(\bar{A}_{t-1}^{\mathsf{d}}),\epsilon_{t})$.
@@ 3.1 Assumptions and sequential regression
To identify the parameter, $\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))]$, the following assumptions, which combine elements of the identification assumptions from D´ıaz et al. (2021) and D´ıaz et al. (2023b), are sufficient. Note that 1 involves future histories rather than single timepoints only.
@@
3Piecewise-smooth invertibility of treatment modification.
@@
Our identification strategy makes use of a representation of longitudinal causal estimands by means of “sequential regression” as proposed by Bang and Robins (2005) and further investigated by van der Laan and Gruber (2012); Luedtke et al. (2017), and Rotnitzky et al. (2017). This approach considers a recursive regression of outcomes (or pseudo-outcomes) onto previous variables, starting at $t=\tau$ and continuing backward to $t=0$. As in D´ıaz et al. (2023b), we split the recursion over the two non-interventional variables $L$ and $Z$ as follows. We consider fixed, pre-specified interventions $\mathsf{d}^{\prime}$, $\mathsf{d}^{\star}$, and $\bar{m}$. Starting with $\mathsf{Q}_{Z,\tau+1}=Y$, define
@@
When the meaning is clear, noting that the only non-random components of the functions $\mathsf{Q}$ are the user-specified $\mathsf{d}^{\prime},\mathsf{d}^{\star}$, and $\bar{m}$, we may use the abbreviations $\mathsf{Q}_{L,t}^{\prime}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{L,t}(\bar{H}_{M,t},\underaccent{\bar}{m}_{t})$, $\mathsf{Q}_{Z,t}^{\prime}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{Z,t}(A^{{}^{\prime}}_{t},\bar{H}_{A,t},\underaccent{\bar}{m}_{t})$, and $\mathsf{Q}_{M,t}^{\star}(\underaccent{\bar}{m}_{t})=\mathsf{Q}_{M,t}(A_{t}^{\star},\bar{H}_{A,t},\underaccent{\bar}{m}_{t})$. We give the following identification result, which is a generalization of Theorem 1 of D´ıaz et al. (2023b).